英語 (English)

発信型英語 (Communicative English) (夜間主 ((化電知)2 年，全 3 ・ 4 年))

ワイスドナルド · 非常勤講師 / 全学共通教育センター 2 単位前期火 13 ・ 14

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】 The purpose of this class is to develop the student’s speaking and

listening skills for everyday English in an international society. 【授業の概要】 TOPICS: Meeting and getting to know people, expressing preferences,

describing and asking about everyday life and life-styles, shopping, eating out, asking about and describing various countries.

【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】

1. Talk about life in Japan and other countries using English.

2. Learn the English you will need to make friends. 【授業の計画】

1. Introduction to the course; self introductions; making introductions.

2. Daily activities: yourself, your country, other countries.

3. Shopping and cost of living in various countries.

4. Entertainment: What do you like? What’s popular in Japan? In other countries?

5. Review lessons 1- 4; Self-assessment and mini-test.

6. Families

7. Sports, games, and exercise.

8. What do you do for fun?

9. Neighborhoods around the world.

10. Review lessons 6- 9; Self assessment and mini-test.

11. Describing people: appearance, behavior, fashion, and style.

12. Wonderful and strange activities all over the world.

13. Tourist destinations: great places to go.

14. 1. Eating out around the world. 2. Preparation for the final exam.

15. Final exam

16. Course review.

【教科書】 TITLE: New Interchange Student Book 1; Author - Jack C. Richards

Publisher: Cambridge University Press 【参考書等】 [参考資料] 【成績評価の方法】 Final exam 50% Mini-tests 20% Class participation and homework

30%. 【再試験の有無】 [再評価] 【受講者へのメッセージ】 Please be prepared to talk about the world in our class.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221354 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ ワイス . (オフィスアワー: woosongdonald@yahoo.com)

English

Communicative English (夜間主 ((化電知)2 年，全 3 ・ 4 年))

Donald Weiss · Part-time Lecturer / Center for General Education 2 units 前期火 13 ・ 14

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ The purpose of this class is to develop the student’s speaking and listening

skills for everyday English in an international society. Outline⟩ TOPICS: Meeting and getting to know people, expressing preferences,

describing and asking about everyday life and life-styles, shopping, eating out, asking about and describing various countries.

�������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩

1. Talk about life in Japan and other countries using English.

2. Learn the English you will need to make friends.

Schedule⟩

1. Introduction to the course; self introductions; making introductions.

2. Daily activities: yourself, your country, other countries.

3. Shopping and cost of living in various countries.

4. Entertainment: What do you like? What’s popular in Japan? In other countries?

5. Review lessons 1- 4; Self-assessment and mini-test.

6. Families

7. Sports, games, and exercise.

8. What do you do for fun?

9. Neighborhoods around the world.

10. Review lessons 6- 9; Self assessment and mini-test.

11. Describing people: appearance, behavior, fashion, and style.

12. Wonderful and strange activities all over the world.

13. Tourist destinations: great places to go.

14. 1. Eating out around the world. 2. Preparation for the final exam.

15. Final exam

16. Course review. Textbook⟩ TITLE: New Interchange Student Book 1; Author - Jack C. Richards

Publisher: Cambridge University Press Reference⟩ [参考資料] Evaluation Criteria⟩ Final exam 50% Mini-tests 20% Class participation and

homework 30%. Re-evaluation⟩ [再評価] Message⟩ Please be prepared to talk about the world in our class.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221354 Contact (Office-Hour, Room, E-mail)⟩

⇒ Weiss . (Office Hour: woosongdonald@yahoo.com)